# revtex4-2Repair the float

### v1

@@ III Interacting regime and Kondo screening
Having obtained the hybridization function $\Gamma(\omega)$, we are ready to use the NRG method to address the Kondo effect in the system. This numerical technique is a powerful and standard impurity solver, which consists of a logarithmic discretization of the continuum conduction band in energy scales that decreases a $\Lambda^{-N/2}$, where $\Lambda$ is a discretization parameter and $N$ is the number of iteration in the NRG procedure. The discretized model is then mapped into a one-dimensional tight-binding-like Hamiltonian (aka Wilson chain Hamiltonian) whose length increase iteratively with $N$. The iterative diagonalization of a large-$N$ Wilson chain is possible by proper truncation of Hilbert space at certain maximum dimension, allowing to access the energy around the Fermi level in a controllable manner [30]. Here, for practical purpose, we employ the NRG method to calculate the impurity retarded Green’s function, from which we calculate the impurity density of states in the interacting regime. Within the Zubarev’s notation [31], our Green’s function can be written as
@@
The nontrivial energy dependence of the hybridization function requires an improved discretization scheme to reproduce with high resolution the initial function, thus reducing the numeric artifacts that can be present in the NRG calculations. To this end, we have used the adaptive “z-averaging” scheme [32] (with $N_{z}$=5), as implemented in the NRG Ljubljana open source code [33]. Furthermore, for the spectral functions calculations we have also used the self-energy “trick”, as described in [34] to remove nonphysical oscillations at the vicinity $\omega=0$. For all the NRG calculations, we have assumed (unless stated otherwise) $\varepsilon_{d}=-U/2$ (particle-hole symmetry), U=0.25, $\Lambda$=2 and number of kept states = 2000.
@@ III.1 Kondo temperature
We start by investigating the effect of the AM exchange field in the Kondo temperature of the system. The main quantity calculated here is the impurity density of state,
@@
It is known that $A(\omega)=\sum_{\sigma}A_{\sigma}(\omega)$ exhibits a many-body resonance at the Fermi level, known as Abrikosov-Shul resonance [35, 36] (or simply Kondo peak), as a signature of the Kondo screening [1]. This resonance not only signals the presence of Kondo screening in the system, but also provides a useful way to estimate $T_{K}$. Indeed, the width of the Kondo resonance can be directly associated to $T_{K}$ [37, 38, 39]. In Fig. 5(a)-5(d) we make an insightful analysis on the behavior of $T_{K}$, extracted from the half-width at half-peak of the impurity spectral function, for different configurations of $t_{AM}$ and $\lambda$, setting ${\bm{B}}=0$. Fig. 5(a) shows $A(\omega)$ for different values of $t_{AM}$. For $t_{AM}=0$ (black line), we see that $A(\omega)$ exhibits a split peak around $\omega=0$. This splitting is induced by the sharp peak in the hybridization function $\Gamma(\omega)$ shown in the Fig. 3(a) for $t_{\rm AM}=0$ (black curve). This splitting is analogous to the one observed in Ref. [40], resulting from an Anderson impurity coupled to a structured host density of states provided by a non-interacting energy level. Here, on the other hand, the sharp structure the conduction band is naturally provided by the van Hove singularity of the two-dimensional material. Now, as $t_{\rm AM}$ increases, we observe that Kondo peak evolves rapidly to a single and sharper peak, as already observed for $t_{\rm AM}=0.05$. This is because the sharp van Hove singularity is split for finite $T_{\rm AM}$, leaving $\Gamma(\omega)$ relatively smooth near the Fermi level, as seen in Fig. 3(a) for finite $t_{\rm AM}$. This prompt sharpening of the Kondo resonance, as $t_{\rm AM}$ increases reveals a strong dependence of $T_{\rm K}$ with $t_{\rm AM}$. Fig. 5(c) shows how $T_{\rm K}$ decreases as $t_{\rm AM}$ increases as anticipated earlier in this paper.
Figure 5: (a) Impurity spectral function $A(\omega)$ for different values of $t_{AM}$ with $\lambda$=0. (b) Impurity spectral function $A(\omega)$ for different values of $\lambda$ with $t_{AM}$=0. (c) $T_{K}$ as function of $t_{AM}$. (d) $T_{K}$ as function of $\lambda$. For all panels, $\bm{B}$ is set to zero.
To compare the effects of $t_{\rm AM}$ to $\lambda$ on $T_{\rm K}$, in Fig. 5(b) shows how $A(\omega)$ is affected the spin-orbit coupling $\lambda$. We now set $t_{\rm AM}=0$. We note that $\lambda$ produces, qualitatively, a very similar to the effect of $t_{\rm AM}$, although much less pronounced, as $\lambda$ produces a very modest splitting of the $\Delta(\omega)$. Indeed, the effect of $t_{\rm AM}$ on $T_{\rm K}$ is much more important as compared to the effect of $\lambda$. By comparing Fig. 5(c) and Fig. 5(d) we observe that $t_{\rm AM}=0.3$ produces a suppression of $t_{\rm K}$ almost three order of magnitude, while for the same value of $\lambda$ the suppression is about one order of magnitude.
@@
Here, we should mention that the dependence of $T_{\rm K}$ with $\lambda$ has been subject of interest over the past decades, and it is still source of debate in the literature [15]. For instance, Zarea et al. reported that away from the particle-hole symmetry in a 2DEG, the parity breaking Dzyaloshinsky-Moriya term is able to renormalize the antiferromagnetic Kondo coupling with an exponential enhancement of $T_{K}$ [13]. A similar result has also been found by one of us for a quantum wire, even at the particle-hole symmetric point [41]. Interestingly, depending on how drastic is the modification of conduction band, we can also observe a suppression of $T_{K}$, as the RSOC is increased, as reported for a quantum wire in the absence of $\bm{B}$ [14].
Within the NRG approach, its is quite clear that the change in the Kondo temperature is directly associated to the modification around the Fermi level in the density of states of the host material induced by properties of the band structure. Therefore, for a given fixed coupling $V$ as in Eq. (4), enhanced density of the host material produces an enhancement of the $T_{\rm K}$. A splitting of the Kondo peak is produced if $\rho(\omega)$ exhibit a peak at $\omega=0$ whose width is similar or smaller than $T_{K}$. See for instance the discussion in Ref. [40].
@@ III.2 Effect of magnetic field
Let us start looking the Fig. 6 (a) where there is no magnetic field applied, therefore all the projected component of $A_{\sigma}$ are equal to each other. Interestingly, even though TRS is broken by AM exchange, because of the unusual spin splitting in momentum space, both spin-components of the $\bm{G}_{\rm host}$ contributes equally to the integral of Eq. 6. As expected, an inversion symmetry breaking induced by RSOC does not split $A_{\sigma}(\omega)$, as shown in Fig. 6 (e). Moreover, the results are almost unaffected when RSOC is included. Indeed, as observed in the insets of Fig. 6(a) and Fig. 6(b), the major splitting in the hybridization function is produced by $t_{\rm AM}$.
Now, intriguing behavior occur when the magnetic field is considered. In Fig. 6(b)-(d) and Fig.6(f)-6(h) show LDOS projected along the different directions of applied $\bm{B}$. As such, for instance, if the magnetic field is applied along the $x$-direction, $\uparrow$ and $\downarrow$ refer to spin pointing towards positive and negating $x$-direction   [^1] To obtain the LDOS projected along different spin directions we calculate the corresponding Green’s function, which can be obtained by combining appropriately the components given by Eq. (7). . Having this in mind, we observe, that regardless the direction of the applied $\bm{B}$, a suppression of the Kondo peak accompanied by splitting of the Kondo peak. However, for an in-plane magnetic field, its interplay with the AM term and with $\bm{B}_{\rm SOC}$ brings a remarkable anisotropic effect in the LDOS as shown in Fig. 6(b)-6(c). Indeed, we see a stronger suppression of $A_{\sigma}(\omega)$ as compared to $\bm{B}$ along the $z$-direction in Fig. 6(d) and (h). This anisotropic effect is observed only if the direct Zeeman effect is negligible in the impurity ($g_{\rm imp}\approx 0$). If we consider $g_{\rm imp}=0.4$ as shown in dashed lines, the Kondo resonance peak is almost fully suppressed, as at this magnetic field strength both spins (band and impurity) are fully polarized, thus completely destroying the Kondo singlet state.
@@
Before concluding, in Fig. 7 we show how the anisotropic magnetic response manifests in the total density of states, $A(\omega)$. To this end, we show $A(\omega)$ for different strengths and direction of the applied magnetic field. Upper and lower panels correspond to $g_{\rm imp}=0$ and $g_{\rm imp}=0.4g$. For $B=0$ the Kondo peak can be clearly seen as shown in black line in Fig 7(a). Now, as $\bm{B}$ increases the Kondo peak is progressively suppressed for magnetic field along either $\bm{\hat{\textbf{x}}}$, $\bm{\hat{\textbf{y}}}$ and $\bm{\hat{\textbf{z}}}$ direction. As shown in Fig. 7(a) and 7(b) for $g_{\rm imp}$=0. However, for field along $z$-direction the peak is less affected, as compared to the other directions. Again, this a signature of the anisotropic magnetic response of the system induced by the interplay between $\bm{B}$, AM and RSOC. Finally, for $g_{\rm imp}=0.4g$, the Kondo peak is strongly suppressed with an enhanced splitting, as shown in Fig. 7 (c)-(d), providing support that the system no longer sustain the Kondo single state.
@@ IV Conclusions
In summary, we have studied the Kondo screening effect in a spin-1/2 impurity coupled to a two-dimensional altermagnet in the presence of external magnetic field. By describing the system within a single-impurity Anderson Hamiltonian, we determined the hybridization function by exact calculation of the local Green’s function, which allows us to garner the complete influence of the host material onto the impurity. We employ a numerical renormalization group approach to calculate the spin-resolved local density of states in the interacting regime, which allowed us to attain the low-temperature Kondo physics in the system. Our results shows a drastic decrease of the Kondo temperature $T_{\rm K}$ as the altermagnet coupling $t_{\rm AM}$ increases. By comparing this suppression with the one produced by the RSOC $\lambda$, we showed that the $t_{\rm AM}$ has a dominant effect in the suppression of $T_{\rm K}$. Furthermore, despite the time-reversal symmetry broken by $t_{\rm AM}$, it does not produce spin-spliting in the hybridization function. This is a direct consequence of the alternating momentum-dependent band spin-splitting that renders zero net magnetization to the impurity upon integration of the self-energy over the entire Brillouin zone. Moreover, we have investigated the effect of an applied magnetic field along different directions. The results showed an important anisotropic response of the system to the applied field. Indeed, a in-plane field has strong effect in Kondo screening marked by a sizable suppression of Kondo peak. In contrast, this suppression is less pronounced for a magnetic field along the $z$-direction. When the Zeeman coupling takes place directly in the impurity, the Kondo peak is overwhelmingly suppressed by the field and the anisotropic effect blurred. Our work shed light on the Kondo correlation in altermagnets and paves the way for future theoretical as well as experimental investigations of correlated phenomena in these materials.
@@ Acknowledgements.
The authors acknowledge financial support from CAPES, FAPEMIG and CNPq. EV thanks FAPEMIG (Process PPM- 00631-17) and CNPq (Process 311366/2021-0). This work used resources of the “Centro Nacional de Processamento de Alto Desempenho em São Paulo (CENAPAD-SP).”

### v2

@@ III Interacting regime and Kondo screening
Having obtained the hybridization function $\Gamma(\omega)$, we are ready to use the NRG method to address the Kondo effect in the system. This numerical technique is a powerful and standard impurity solver, which consists of a logarithmic discretization of the continuum conduction band in energy scales that decreases a $\Lambda^{-N/2}$, where $\Lambda$ is a discretization parameter and $N$ is the number of iterations in the NRG procedure. The discretized model is then mapped into a one-dimensional tight-binding-like Hamiltonian (aka Wilson chain Hamiltonian) whose length increase iteratively with $N$. The iterative diagonalization of a large-$N$ Wilson chain is possible by proper truncation of Hilbert space at certain maximum dimension, allowing us to access the energy around the Fermi level in a controllable manner [34]. Here, for practical purposes, we employ the NRG method to calculate the impurity retarded Green’s function, from which we calculate the impurity density of states in the interacting regime. Within the Zubarev’s notation [35], our Green’s function can be written as
@@
The nontrivial energy dependence of the hybridization function requires an improved discretization scheme to reproduce with high resolution the initial function, thus reducing the numeric artifacts that can be present in the NRG calculations. To this end, we have used the adaptive $z$-averaging scheme [36] (with $N_{z}$=5), as implemented in the NRG Ljubljana open source code [37]. Furthermore, for the spectral functions calculations we have also checked with the self-energy trick, as described in Ref. [38] to analyze and remove nonphysical oscillations at the vicinity $\omega=0$. For all the NRG calculations, we have assumed (unless stated otherwise) $\varepsilon_{d}=-U/2$ (particle-hole symmetry), U=0.25, $\Lambda$=2, and number of kept states = 2000.
@@ III.1 Kondo temperature
We start by investigating the effect of the AM exchange field in the Kondo temperature of the system. The main quantity calculated here is the impurity density of states [39]:
@@
It is known that $A(\omega)=\sum_{\sigma}A_{\sigma}(\omega)$ exhibits a many-body resonance at the Fermi level, known as Abrikosov-Shul resonance [40, 41] (or simply Kondo peak), as a signature of the Kondo screening [1]. This resonance not only signals the presence of Kondo screening in the system, but also provides a useful way to estimate $T_{K}$. Indeed, the width of the Kondo resonance can be directly associated to $T_{K}$ [42, 43, 44]. In Fig. 5(a)-5(d), we make an insightful analysis on the behavior of $T_{K}$, extracted from the half-width at half-peak of the impurity spectral function, for different configurations of $t_{\rm AM}$ and $\lambda$, setting ${\bm{B}}=0$. Fig. 5(a) shows $A(\omega)$ for different values of $t_{\rm AM}$. For $t_{\rm AM}=0$ (black line), we see that $A(\omega)$ exhibits a split peak around $\omega=0$. This splitting is induced by the sharp peak in the hybridization function $\Gamma(\omega)$ shown in Fig. 3(a) for $t_{\rm AM}=0$ (black curve). This splitting is analogous to the one observed in Ref. [45], resulting from an Anderson impurity coupled to a structured host density of states provided by a non-interacting energy level. Here, on the other hand, the sharp structure the conduction band is naturally provided by the van Hove singularity of the two-dimensional material. Now, as $t_{\rm AM}$ increases, we observe that the Kondo peak evolves rapidly to a single and sharper peak, as already observed for $t_{\rm AM}=0.05$. This is because the sharp van Hove singularity is split for finite $T_{\rm AM}$, leaving $\Gamma(\omega)$ relatively smooth near the Fermi level, as seen in Fig. 3(a) for finite $t_{\rm AM}$. This prompt sharpening of the Kondo resonance, as $t_{\rm AM}$ increases reveals a strong dependence of $T_{\rm K}$ with $t_{\rm AM}$. Fig. 5(c) shows how $T_{\rm K}$ decreases as $t_{\rm AM}$ increases as anticipated earlier in this paper.
Figure 5: (a) Impurity spectral function $A(\omega)$ for different values of $t_{\rm AM}$ with $\lambda$=0. (b) Impurity spectral function $A(\omega)$ for different values of $\lambda$ with $t_{\rm AM}$=0. (c) $T_{K}$ as function of $t_{\rm AM}$. (d) $T_{K}$ as function of $\lambda$. For all panels, $\bm{B}$ is set to zero. The insets in Fig. 5 (a)-(b) are zooms for the range of $\omega$ in [-0.01,0.01] for better visualization.
To compare the effects of $t_{\rm AM}$ to $\lambda$ on $T_{\rm K}$, in Fig. 5(b) shows how $A(\omega)$ is affected by the SOC $\lambda$. We now set $t_{\rm AM}=0$. We note that $\lambda$ produces, qualitatively, a very similar effect of $t_{\rm AM}$, although much less pronounced, as $\lambda$ produces a very modest splitting of the $\Delta(\omega)$. Indeed, the effect of $t_{\rm AM}$ on $T_{\rm K}$ is much more important as compared to the effect of $\lambda$. By comparing Figs. 5(c) and 5(d), we observe that $t_{\rm AM}=0.3$ produces a suppression of $T_{\rm K}$ by almost three orders of magnitude, while for the same value of $\lambda$ the suppression is about one order of magnitude.
@@
Here, we should mention that the dependence of $T_{\rm K}$ with $\lambda$ has been a subject of interest over the past decades, and it is still source of debate in the literature [15]. For instance, Zarea et al. [13] reported that away from the particle-hole symmetry in a two-dimensional electron gas (2DEG), the parity breaking Dzyaloshinsky-Moriya term is able to renormalize the antiferromagnetic Kondo coupling with an exponential enhancement of $T_{K}$ [13]. A similar result has also been found by one of us for a quantum wire, even at the particle-hole symmetric point [46]. Interestingly, depending on how drastic the modification of conduction band is, we can also observe a suppression of $T_{K}$, as the RSOC is increased, as reported for a quantum wire in the absence of $\bm{B}$ [14].
Within the NRG approach, its is quite clear that the change in the Kondo temperature is directly associated to the modification around the Fermi level in the density of states of the host material induced by properties of the band structure. Therefore, for a given fixed coupling $V$ as in Eq. (4), enhanced density of the host material produces an enhancement of the $T_{\rm K}$. A splitting of the Kondo peak is produced if the host density of state, $\rho(\omega)$ exhibit a peak at $\omega=0$ whose width is similar or smaller than $T_{K}$. See, for instance, the discussion in Ref. [45].
@@ III.2 Effect of magnetic field
Let us start looking at Fig. 6 (a) where there is no magnetic field applied, therefore all the projected components of $A_{\sigma}$ are equal to each other. Interestingly, even though TRS is broken by AM exchange, because of the unusual spin splitting in momentum space, both spin-components of the $\bm{G}_{\rm host}$ contributes equally to the integral of Eq. (6). As expected, an inversion symmetry breaking induced by RSOC does not split $A_{\sigma}(\omega)$, as shown in Fig. 6 (e). Moreover, the results are almost unaffected when RSOC is included. Indeed, as observed in the insets of Figs. 6(a) and 6(b), the major splitting in the hybridization function is produced by $t_{\rm AM}$.
Now, intriguing behavior occur when the magnetic field is considered. Figures  6(b)-6(d) and 6(f)-6(h) show LDOS projected along the different directions of applied $\bm{B}$. As such, for instance, if the magnetic field is applied along the $x$-direction, $\uparrow$ and $\downarrow$ refer to spin pointing towards positive and negative $x$-direction [47]. Keeping this in mind, we observe, that regardless of the direction of the applied $\bm{B}$, a suppression of the Kondo peak is accompanied by splitting of the Kondo peak. However, for an in-plane magnetic field, its interplay with the AM term and with $\bm{B}_{\rm SOC}$ brings a remarkable anisotropic effect in the LDOS as shown in Fig. 6(b) and 6(c). Indeed, we see a stronger suppression of $A_{\sigma}(\omega)$ as compared to $\bm{B}$ along the $z$-direction in Fig. 6(d) and 6(h). This anisotropic effect is observed only if the direct Zeeman effect is negligible in the impurity ($g_{\rm imp}\approx 0$). If we consider $g_{\rm imp}=0.4$ as shown in dashed lines, the Kondo resonance peak is almost fully suppressed, as at this magnetic field strength both spins (band and impurity) are fully polarized, thus completely destroying the Kondo singlet state.
@@
Before concluding, in Fig. 7 we show how the anisotropic magnetic response manifests in the total density of states, $A(\omega)$. To this end, we show $A(\omega)$ for different strengths and direction of the applied magnetic field. Upper and lower panels correspond to $g_{\rm imp}=0$ and $g_{\rm imp}=0.4g$. For $B=0$, the Kondo peak can be clearly seen as shown by the black line in Fig 7(a). Now, as $\bm{B}$ increases the Kondo peak is progressively suppressed for magnetic field along either the $\bm{\hat{\textbf{x}}}$, $\bm{\hat{\textbf{y}}}$ or $\bm{\hat{\textbf{z}}}$ direction, as shown in Fig. 7(a) and 7(b) for $g_{\rm imp}$=0. However, for the field along $z$-direction the peak is less affected, as compared to the other directions. Again, this is a signature of the anisotropic magnetic response of the system induced by the interplay between $\bm{B}$, AM, and RSOC. Finally, for $g_{\rm imp}=0.4g$, the Kondo peak is strongly suppressed with an enhanced splitting, as shown in Fig. 7 (c) and 7(d), providing support that the system no longer sustains the Kondo single state.
@@ IV Conclusions
In summary, we have studied the Kondo screening effect in a spin-1/2 impurity coupled to a two-dimensional altermagnet in the presence of external magnetic field. By describing the system within a single-impurity Anderson Hamiltonian, we determined the hybridization function by exact calculation of the local Green’s function, which allows us to garner the complete influence of the host material onto the impurity. We employ a NRG approach to calculate the spin-resolved local density of states in the interacting regime, which allowed us to attain the low-temperature Kondo physics in the system. Our results shows a drastic decrease of the Kondo temperature $T_{\rm K}$ as the altermagnet coupling $t_{\rm AM}$ increases. By comparing this suppression with the one produced by the RSOC $\lambda$, we showed that $t_{\rm AM}$ has a dominant effect in the suppression of $T_{\rm K}$. Furthermore, despite the TRS broken by $t_{\rm AM}$, it does not produce spin-splitting in the hybridization function. This is a direct consequence of the alternating momentum-dependent band spin-splitting that renders zero net magnetization to the impurity upon integration of the self-energy over the entire Brillouin zone. Moreover, we have investigated the effect of an applied magnetic field along different directions. The results showed an important anisotropic response of the system to the applied field. Indeed, a in-plane field has a strong effect on Kondo screening marked by a sizable suppression of the Kondo peak. In contrast, this suppression is less pronounced for a magnetic field along the $z$-direction. When the Zeeman coupling takes place directly in the impurity, the Kondo peak is overwhelmingly suppressed by the field and the anisotropic effect blurred. Our paper shed light on the Kondo correlation in altermagnets and paves the way for future theoretical as well as experimental investigations of correlated phenomena in these materials.
@@ Acknowledgements.
The authors acknowledge financial support from CAPES, FAPEMIG and CNPq. EV thanks FAPEMIG (Process No. PPM- 00631-17) and CNPq (Process No. 311366/2021-0). This paper used resources of the “Centro Nacional de Processamento de Alto Desempenho em São Paulo (CENAPAD-SP).”